# GeoSplat: A Deep Dive into Geometry-Constrained Gaussian Splatting

### v2

@@ Appendix B More Preliminaries on Manifold: Unique Connection
The affine connection $\nabla^{\mathcal{M}}$ that is compatible with the metric $g^{\mathcal{M}}$ turns out to be unique. The Koszul identity (Lee, 2018) provides an explicit formula for this connection. We show its derivation process as below. Based on the metric compatibility, we have
@@ Appendix D Adapted Expression of Varifold-based Curvatures
Buet et al. (2022) provided an analytical solution of weak second fundamental form (WSFF) for varifold. The aim of this section is to adapt that solution to our setting.
@@
To obtain full curvature information from the varifold $\mathcal{W}$, Buet et al. (2022) first introduced a type of variation called G-linear $\vartheta_{i,j,k}\mathcal{W}$ as
$\vartheta_{i,j,k}\mathcal{W}(\varphi)=\int_{\Omega\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}}\mathsf{T}_{j,k}\langle\nabla^{\mathsf{T}}\varphi(\mathbf{q}),\mathbf{e}_{i}\rangle\mathrm{d}\mathcal{W}(\mathbf{q},\mathsf{T}),$ (29)
where $\varphi:\mathsf{M}\rightarrow\mathbb{R}$ is an arbitrary differentiable function, $\nabla^{\mathsf{T}}$ represents the gradient clipped by tangential projection matrix $\mathsf{T}$, and $\mathbf{e}_{i}$ denotes the $i$-th Euclidean basis vector. Similar to the first-order variation $\vartheta\mathcal{W}$, this new one also admits a decomposition as
$\vartheta_{i,j,k}\mathcal{W}=-b_{i,j,k}\|\mathcal{W}\|+\widetilde{\vartheta}_{i,j,k}\mathcal{W},$ (30)
if it is a Radon measure for any $i,j,k\in[1,D_{\mathcal{E}}]$. Here $b_{i,j,k}:\mathsf{M}\rightarrow\mathbb{R}$ is a continuous function, and $\widetilde{\vartheta}_{i,j,k}\mathcal{W}$ is a measure singular with respect to $\|\mathcal{W}\|$. Then, in the case where $\mathsf{W}=\{\mathbf{q}_{i}\}_{i\in\mathcal{I}}$ is a discrete point set and every point $\mathbf{q}_{i}$ is paired with a tangent plane $\mathsf{T}_{i}$, Buet et al. (2022) derived a closed-form WSFF based on the coefficient $b_{i,j,k}$ as
@@
where $m_{l_{2}},l_{2}\in\mathcal{I}$ is the point weight and $\Upsilon,\chi$ are two kernel functions. The curvature quantities can be finally obtained through the eigendecomposition of the below matrix:
@@
where :: means to take all possible entries to form a vector, $\widebar{\mathsf{T}}_{l_{1}}$ denotes an orthonormal basis of the tangent space $\mathsf{T}_{l_{1}}$, $\mathbf{n}_{l_{1}}$ is a normal vector perpendicular to the space.
@@
Now, let us adapt the above analytical WSFF $\mathcal{B}$ to our setting, facilitating practical computation. As the first step, we specify the rectifiable set as the positions of Gaussian primitives: $\mathsf{M}=\{\bm{\mu}^{i}_{\mathrm{obj}}\}_{i\in\mathcal{I}}$, with each element $\bm{\mu}^{i}_{\mathrm{obj}}$ paired with a tangent plane $\widetilde{\mathbf{V}}^{i}$ that is determined by the normal vector $\widetilde{\mathbf{u}}^{i}_{3}$ as $\mathbf{I}-\widetilde{\mathbf{u}}^{i}_{3}(\widetilde{\mathbf{u}}^{i}_{3})^{\top}$. Here $D_{\mathcal{M}}=2$ for novel view rendering in the 3D world $\mathcal{E}=\mathbb{R}^{3}$. For better notational consistencies, we also relabel the middle term $\mathbf{B}_{i,j,k}(\mathbf{q}_{l_{1}})$ as $\mathsf{B}_{d_{1},d_{2},d_{3}}^{i}$ (given $\mathbf{q}_{l_{1}}=\bm{\mu}^{i}_{\mathrm{obj}}$) and the subscript $l_{2}$ as $j$.
@@
As the tangential projection of a normal vector is zero: $\widetilde{\mathbf{V}}^{i}\widetilde{\mathbf{u}}^{i}_{3}=\mathbf{0}$, we have
@@
given that $D_{\mathcal{M}}=2,D_{\mathcal{E}}=3$. Therefore, we express WSFF as an interpolated matrix:
@@
where $\mathrm{sym}(\mathbf{A})$ is an operation $\frac{1}{2}(\mathbf{A}+\mathbf{A}^{\top})$ defined for any matrix.
To summarize, the adapted WSFF for our setting is as
@@
where we relabel the origional kernel functions as $\chi_{\epsilon},\Upsilon^{\prime}_{\epsilon}$ that depends on bandwidth $\epsilon$.
Appendix E Manifold-based Geometric Estimations
@@ E.1.2 Analysis of the Tangent Space Structure
The proof is provided in Appendix F. ∎
@@
The proof is provided in Appendix G. ∎
@@ E.1.3 Laplacian Operator Estimation
The proof is provided in Appendix H. ∎
@@
$\left\{\begin{aligned} &k_{t}(\mathbf{q}_{1},\mathbf{q}_{2})\propto\exp\Big((\mathbf{q}_{1}-\mathbf{q}_{2})^{\top}\bm{\Lambda}_{t}(\mathbf{q}_{1},\mathbf{q}_{2})^{-1}(\mathbf{q}_{1}-\mathbf{q}_{2})\Big)\\ &\bm{\Lambda}_{t}(\mathbf{q}_{1}=\mathbf{c}^{i}_{\mathrm{obj}},\mathbf{q}_{2}=\mathbf{c}^{j}_{\mathrm{obj}})=t\mathbf{\Sigma}^{i}_{\mathrm{obj}}\mathbf{\Sigma}^{j}_{\mathrm{obj}}\end{aligned}\right.$. (47)
@@ E.2 Dimension and Normal Vector Estimations
where $\times$ denotes the vector cross product in three dimensions ($\mathcal{D}_{\mathcal{E}}=3$). More conveniently, both manifold dimension $D_{\mathcal{M}}$ and normal space $\mathcal{N}_{\mathbf{q}}\mathcal{M}$ are just side products from the kernel matrix decomposition in Theorem E.2. The details are as below.
@@
The proof is provided in Appendix I. ∎
@@ E.3 Curvature Estimation
The proof is provided in Appendix J. ∎
@@
The proof is provided in Appendix K. ∎
@@
Appendix F Proof: Estimation of the Tangent Space
@@ F.2 Rank Determination
F.3 Kernel Matrix Analysis
@@
Appendix G Proof: Curve-based Formulation of the Riemannian Metric
@@
G.1 Analysis on the Euclidean Space
@@
G.2 Analysis on the Curved Manifold
@@
Appendix H Proof: Markovian Characterization of the Linear Operator
@@
Appendix I Proof: Estimations of the Dimension and Normal Vectors
For an arbitrary point $\mathbf{q}=\bm{\mu}^{i}_{\mathrm{obj}}$ residing in the embedded Riemannian manifold $\mathcal{M}\hookrightarrow\mathcal{E}$, we have proved two key conclusions in Appendix F.3:
@@ Appendix I Proof: Estimations of the Dimension and Normal Vectors
Appendix J Proof: Curved-based Formulation of the Shape Operator
@@
Appendix K Proof: Curvature Estimation
@@
Appendix L Experiment Settings and More Results
@@
The vanilla Gaussian splatting (i.e., 3DGS) (Kerbl et al., 2023) is a natural baseline in our setting for verifying the performance gains of our framework. In addition to this, an important baseline is GeoGaussian (Li et al., 2025), which adopted a number of regularization strategies (e.g., co-planar constraints) based on low-order geometric priors (e.g., normal information) and covered most of other works (Wang et al., 2024; Turkulainen et al., 2025) on geometric regularization. It is necessary to show that our framework, which is compatible with their low-order regularization and further exploits high-order geometric information, can achieve better performance. We also compare with Ververas et al. (2025) to show that the mean absolute curvature (MAC) is indeed a better low-curvature area identifier than the mean curvature.
@@
L.2 More Case Studies
@@
L.3 Enriched Gaussian Primitives

### v3

@@ Appendix B More Preliminaries on Manifold: Unique Connection
The affine connection $\nabla^{\mathcal{M}}$ that is compatible with the metric $g^{\mathcal{M}}$ turns out to be unique. The Koszul identity (Lee, 2018) provides an explicit formula for this connection. We show below its derivation process. Based on the metric compatibility, we have
@@ Appendix D Adapted Expression of Varifold-based Curvatures
Buet et al. (2022) introduced a notion of approximate second fundamental form (WSFF) for point cloud varifolds, with convergence and consistency properties. The aim of this section is to adapt that notion to our setting.
@@
To obtain full curvature information from a varifold $\mathcal{W}$, Buet et al. (2022) first introduced a collection of G-linear variations indexed by $i,j,k$ and defined by their action on $C_{c}^{1}(\Omega)$ functions as
$\delta_{i,j,k}\mathcal{W}(\varphi)=\int_{\Omega\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}}\mathsf{T}_{j,k}\langle\nabla^{\mathsf{T}}\varphi(\mathbf{q}),\mathbf{e}_{i}\rangle\mathrm{d}\mathcal{W}(\mathbf{q},\mathsf{T}),$ (29)
where every element $T$ of the Grassmannian is identified with its orthogonal projection matrix $(T_{jk})$, $\nabla^{\mathsf{T}}$ represents the tangential gradient with respect to $T$, and $\mathbf{e}_{i}$ denotes the $i$-th Euclidean basis vector. We say that $\mathcal{W}$ has locally bounded variations whenever all G-linear variations are Radon measures, in which case they can be decomposed as
$\delta_{i,j,k}\mathcal{W}=-b_{i,j,k}\|\mathcal{W}\|+{\delta}^{S}_{i,j,k}\mathcal{W}.$ (30)
where $b_{i,j,k}\in L^{1}(\|\mathcal{W}\|)$ and ${\delta}^{S}_{i,j,k}\mathcal{W}$ are Radon measures singular with respect to $\|\mathcal{W}\|$. When $\mathsf{W}$ is a point cloud varifold associated with (point, tangent plane, mass) triplets $\{(\mathbf{q}_{i},\mathsf{T}_{i},m_{i})\}_{i\in\mathcal{I}}$, Buet et al. (2022) introduced a consistent notion of approximate weak second fundamental form (WSFF) defined as
@@
where $\Upsilon,\chi$ are suitable kernel functions, and $\epsilon>0$ is the approximation scale. The approximate curvature quantities can be finally obtained through the eigendecomposition of the matrix
@@
where $i,j,:$ means taking all entries along the last index to form a vector, $\widebar{\mathsf{T}}_{l_{1}}$ denotes an orthonormal basis of the tangent space $\mathsf{T}_{l_{1}}$, and $\mathbf{n}_{l_{1}}$ is a normal vector.
@@
Now, let us adapt the above approximate WSFF $\mathcal{B}$ to our setting, facilitating practical computation. As the first step, we specify the points as the positions of Gaussian primitives, $\mathsf{M}=\{\bm{\mu}^{i}_{\mathrm{obj}}\}_{i\in\mathcal{I}}$, with each element $\bm{\mu}^{i}_{\mathrm{obj}}$ paired with a tangent plane $\widetilde{\mathbf{V}}^{i}$ that is determined by the normal vector $\widetilde{\mathbf{u}}^{i}_{3}$ as $\mathbf{I}-\widetilde{\mathbf{u}}^{i}_{3}(\widetilde{\mathbf{u}}^{i}_{3})^{\top}$. Here $D_{\mathcal{M}}=2$ for novel view rendering in the 3D world $\mathcal{E}=\mathbb{R}^{3}$. For better notational consistencies, we also relabel the middle term $\mathbf{B}_{i,j,k}(\mathbf{q}_{l_{1}})$ as $\mathsf{B}_{d_{1},d_{2},d_{3}}^{i}$ (given $\mathbf{q}_{l_{1}}=\bm{\mu}^{i}_{\mathrm{obj}}$) and the subscript $l_{2}$ as $j$.
@@
As the tangential projection of a normal vector is zero, $\widetilde{\mathbf{V}}^{i}\widetilde{\mathbf{u}}^{i}_{3}=\mathbf{0}$, we have
@@
given that $D_{\mathcal{M}}=2,D_{\mathcal{E}}=3$. Therefore, we express the approximate WSFF as an interpolated matrix
@@
where $\mathrm{sym}(\mathbf{A})$ denotes for any matrix $A$ the operation $\frac{1}{2}(\mathbf{A}+\mathbf{A}^{\top})$.
To summarize, the approximate WSFF adapted for our setting is
@@
where we relabel the original kernel functions as $\chi_{\epsilon},\Upsilon^{\prime}_{\epsilon}$ that depend on the approximation scale $\epsilon$.
Appendix E Manifold Structure Estimation for Gaussian Splatting
@@
There are several types of classical frameworks for estimating manifold structures from discrete data. For example, moving least squares (Lancaster & Salkauskas, 1981) and heat equations (van den Berg & Le Gall, 1994; Coifman & Lafon, 2006). Following the direction of heat equations, we provide a full derivation procedure that shows how to practically estimate the manifold-related geometric quantities mentioned in Sec. 2.2, considering both the positions and shapes of Gaussian primitives in our problem setting. While this part should have overlaps with some previous works in this classical direction, our main goal here is to make our method self-contained, providing practical calculation guidance for future research in computer vision and graphics.
@@ E.1.2 Analysis of the Tangent Space Structure
The proof is provided in Appendix E.4. ∎
@@
The proof is provided in Appendix E.5. ∎
@@ E.1.3 Laplacian Operator Estimation
The proof is provided in Appendix E.6. ∎
@@
$\left\{\begin{aligned} &k_{t}(\mathbf{q}_{1},\mathbf{q}_{2})\propto\exp\Big(-(\mathbf{q}_{1}-\mathbf{q}_{2})^{\top}\bm{\Lambda}_{t}(\mathbf{q}_{1},\mathbf{q}_{2})^{-1}(\mathbf{q}_{1}-\mathbf{q}_{2})\Big)\\ &\bm{\Lambda}_{t}(\mathbf{q}_{1}=\mathbf{c}^{i}_{\mathrm{obj}},\mathbf{q}_{2}=\mathbf{c}^{j}_{\mathrm{obj}})=t\mathbf{\Sigma}^{i}_{\mathrm{obj}}\mathbf{\Sigma}^{j}_{\mathrm{obj}}\end{aligned}\right.$. (47)
@@ E.2 Dimension and Normal Vector Estimations
where $\times$ denotes the vector cross product in three dimensions ($\mathcal{D}_{\mathcal{E}}=3$). More conveniently, both manifold dimension $D_{\mathcal{M}}$ and normal space $\mathcal{N}_{\mathbf{q}}\mathcal{M}$ are just side products from the kernel matrix decomposition in Theorem E.2. The details are below.
@@
The proof is provided in Appendix E.7. ∎
@@ E.3 Curvature Estimation
The proof is provided in Appendix E.8. ∎
@@
The proof is provided in Appendix E.9. ∎
@@
E.4 Proof: Estimation of the Tangent Space
@@ E.4.2 Rank Determination
E.4.3 Kernel Matrix Analysis
@@
E.5 Proof: Curve-based Formulation of the Riemannian Metric
@@
E.5.1 Analysis on the Euclidean Space
@@
E.5.2 Analysis on the Curved Manifold
@@
E.6 Proof: Markovian Characterization of the Linear Operator
@@
E.7 Proof: Estimations of the Dimension and Normal Vectors
For an arbitrary point $\mathbf{q}=\bm{\mu}^{i}_{\mathrm{obj}}$ residing in the embedded Riemannian manifold $\mathcal{M}\hookrightarrow\mathcal{E}$, we have proved two key conclusions in Appendix E.4.3:
@@ E.7 Proof: Estimations of the Dimension and Normal Vectors
E.8 Proof: Curved-based Formulation of the Shape Operator
@@
E.9 Proof: Curvature Estimation
@@
Appendix F Experiment Settings and More Results
@@ F.1 Experiment Settings
The vanilla Gaussian splatting (i.e., 3DGS) (Kerbl et al., 2023) is a natural baseline in our setting for verifying the performance gains of our framework. In addition to this, an important baseline is GeoGaussian (Li et al., 2025), which adopted a number of regularization strategies (e.g., co-planar constraints) based on low-order geometric priors (e.g., normal information) and covered most of other works (Wang et al., 2024; Turkulainen et al., 2025) on geometric regularization. It is necessary to show that our framework, which is compatible with their low-order regularization and further exploits higher-order geometric information, can achieve better performance. We also compare with Ververas et al. (2025) to show that the mean absolute curvature (MAC) is indeed a better low-curvature area identifier than the mean curvature.
@@
F.2 More Case Studies
@@
F.3 Enriched Gaussian Primitives